calculate the [OH-] of each aqueous solution with the following [H3O+]:

1: urine, 4.2 • 10^-6 M
2: bile, 7.9 • 10^-9 M
Answer #1

матката Hoot = 4.27100m PH. HoЯ Сн, с*1 — 4 Lч.2110 ) 1 s-37 P+= 1 + 2 + 1-31 por = 8.62 Сонг) 6-Р° -3- - 10. o Сон] = 2.4